# Inflationary relics from an Ultra-Slow-Roll plateau

### v1

@@ 1 Introduction
A particularly well-studied mechanism to enhance the amplitude of curvature perturbations during inflation involves a transient ultra-slow-roll (USR) phase [PhysRevD.50.7173, Tsamis:2003px, Kinney:2005vj]. This regime occurs when the inflaton field encounters a nearly flat region of the potential, causing its velocity to decrease rapidly and the curvature perturbation to grow on super-horizon scales. As a result, the primordial power spectrum can be amplified by several orders of magnitude over a narrow range of scales, leading to the collapse of overdense regions into PBHs after horizon re-entry in the radiation era [Garcia-Bellido:2017mdw, Motohashi:2017kbs, Germani:2017bcs, Byrnes:2018txb, Atal:2018neu, Figueroa:2021zah, Ragavendra:2023ret, Ballesteros:2024zdp, Ballesteros:2024pbe, Fujita:2025imc]. Such an enhancement of the curvature perturbation generically leads to non-Gaussianities (see [Pi:2024lsu] for a review), which can significantly alter the critical conditions for black hole formation, as shown by relativistic numerical studies [2020JCAP...05..022A, 2022JCAP...05..012E, Shimada:2024eec, Inui:2024fgk].
@@
The value of $\beta$ is related to the second derivative of the potential at the top of the small barrier. It was also realized that the relation is not invertible when the linear perturbation $\zeta_{G}$ exceeds $\beta^{-1}$. However, the probability for this to happen is finite, which hinted to a different channel for PBH formation. Indeed, while the barrier slows down the inflaton—enhancing curvature perturbations—it may also trap regions of the field that fluctuate backward, preventing them from crossing the barrier for $\zeta_{G}\gtrsim\beta^{-1}$. These trapped domains evolve into vacuum bubbles, providing an alternative channel of PBH formation distinct from the usual collapse of large adiabatic perturbations [2019JCAP...09..073A, 2020JCAP...05..022A, Escriva:2023uko]. Bubbles in this case form when the background field approaches the top of the barrier, leading to a sharply peaked distribution of bubble sizes. The precise shape of this distribution was determined through dedicated numerical simulations [Escriva:2023uko], which found a scaling behavior for the bubble size. It was also shown that vacuum bubbles (surrounded by type-II fluctuations) can contribute higher PBH abundance than those forming from adiabatic perturbations, if the non-Gaussianity is large, i.e. $\beta\gtrsim 3.1$.
@@ 3 Statistics of initial conditions and shape dispersion
in $U_{+}$ and $U_{-}$ respectively. Here we have introduced $\sigma_{+}^{2}=\int_{U_{+}}{\cal P}_{\delta\phi}\ d\ln k\equiv f\sigma_{\delta\phi}^{2}$, and the complementary $\sigma_{-}^{2}=(1-f)\sigma_{\delta\phi}^{2}$. This leads to
$\Delta W={\mu_{\Delta}^{2}\over\sigma_{\delta\phi}^{2}}{1\over f(1-f)}.$
For fixed $\mu_{\Delta}$ the cost $\Delta W$ is minimized for $f=1/2$, when both subsets in the partition share equal weight in the variance of the field, i.e. $\sigma_{+}^{2}=\sigma_{-}^{2}$. In such case we have $\Delta W=4\mu_{\Delta}^{2}\sigma_{\delta\phi}^{-2}$. Then, from (3.13) the $n=1$ deviation from the mean corresponds to $\mu_{\Delta}=(1/2)\sigma_{\delta\phi}$. It follows from (3.10) and (3.14) that $|\Delta_{\phi}(r)|\leq\mu_{\Delta}$, and therefore we have
@@
Finally, to account for the likelihood of the shapes deviating from the mean profiles, we shall weight those realizations by the factor $f(n,m)=\textrm{Erfc}(\mid n\mid/\sqrt{2})\textrm{Erfc}(\mid m\mid/\sqrt{2})$. An example of the shapes can be found in Fig.5. There, we also display the anticorrelation between the two-point function $\Psi_{\phi}/\Psi_{\pi}$.
@@ 5 Quantum diffusion
[…]
@@ 6 Adiabatic fluctuations and generalized $\zeta[\zeta_{G}]$
$\zeta[\zeta_{G}]=\ln 2-\ln\left(A\,\zeta_{G}+g^{1/3}_{+}+g^{1/3}_{-}\right),$ (6.17)
@@
$g_{\pm}(\zeta_{G})=4-6(2+A)\zeta_{G}+A^{3}\zeta_{G}^{3}\pm 2\sqrt{f},$ (6.18)
$f(\zeta_{G})=(2-3(2+A)\zeta_{G})(2-3(2+A)\zeta_{G}+A^{3}\zeta_{G}^{3}).$ (6.19)
@@
However, from Fig. 9 the value of $A\approx-0.1$ near the peak of the distribution is significant, and cannot be completely neglected.
@@
For each curvature profile $\zeta(r)$ generated with an amplitude $\mu<\mu^{\rm bub}_{c}$ (which therefore does not lead to the formation of vacuum bubbles), we perform relativistic numerical simulations in spherical symmetry to track the collapse of adiabatic curvature fluctuations, from super-horizon scales up to the moment of PBH formation. In particular, we use the numerical code SPriBHoS-II [Escriva:2025eqc] (see also [escriva_solo]), which employs a new methodology based on the Misner–Sharp formalism [PhysRev.136.B571] to follow the hydrodynamic collapse of generic curvature fluctuations (type-I and type-II) by using the trace of the extrinsic curvature. We refer the reader to that work for further details on the numerical procedure. We then iterate over different realizations of $\mu$ to determine the critical threshold conditions, defining this value as $\mu^{\rm adi}_{c}$. This can be converted into a threshold associated with the Gaussian curvature fluctuation $\zeta_{G}$, denoted by $\mu^{\zeta_{G}}_{c}=\mu^{\rm adi}_{c}/\dot{\phi}_{\rm bkg}(N_{*})$, from which we define the peak height $\nu^{\rm adi}_{c}=\mu^{\zeta_{G}}_{c}/\sigma_{\zeta_{G}}(N_{*})$, being $\sigma_{\zeta_{G}}(N_{*})$ the width of the power spectrum of $\mathcal{P}_{\zeta_{G}}(N_{*},k)$. The corresponding numerical results for these cases are shown in Table 1. For the mean profile $(n,m)=(0,0)$ we find that $\mu^{\rm adi}_{c}\approx 2.455\cdot 10^{-5}$ and $\mathcal{C}_{c}(r_{m})\approx 0.485$ (being $r_{m}$ the location of the maximum of the critical compaction function), and we also list the threshold for the other realizations $(n,m)$. When comparing the threshold obtained for the mean profile with that of the analytical template using the case A=0.1A=-0.1 (motivated by Fig.9, which is related to the maximum of the power spectrum), we find good agreement with the numerical result from the $\delta N$, with a deviation in the threshold of about $\sim 3\%$, even the significant deviation when approaching the divergent point discussed previously.
@@
$\beta_{1}(n,m)$ $=\frac{\mathcal{I}(\nu^{\rm bub}_{c}(n,m),\infty)}{\mathcal{I}(\nu^{\rm bub}_{c}(0,0),\infty)}\frac{f(n,m)}{f(0,0)},$ (6.23)
$\beta_{2}(n,m)$ $=\frac{\mathcal{I}(\nu^{\rm adi}_{c}(n,m),\nu^{\rm bub}_{c}(n,m))}{\mathcal{I}(\nu^{\rm adi}_{c}(0,0),\nu^{\rm bub}_{c}(0,0))}\frac{f(n,m)}{f(0,0)},$ (6.24)
@@ Acknowledgments
A.E. thanks the support from the YLC program at the Institute for Advanced Research, Nagoya University. This work is supported in part by the National Key Research and Development Program of China Grant No. 2021YFC2203004. JG is supported by MCIN/AEI/PID2022-136224NB-C22, by the “Center of Excellence Maria de Maeztu 2025-2029” grant CEX2024-001451-M, funded by MICIU/AEI/10.13039/501100011033, and by 2021-SGR00872 funded by AGAUR. S.P. is supported by the National Natural Science Foundation of China (NSFC) Grants Nos. 12475066 and 12447101, and JSPS KAKENHI grant No. 24K00624.

### v2

@@ 1 Introduction
A particularly well-studied mechanism to enhance the amplitude of curvature perturbations during inflation involves a transient ultra-slow-roll (USR) phase [PhysRevD.50.7173, Tsamis:2003px, Kinney:2005vj]. This regime occurs when the inflaton field encounters a nearly flat region of the potential, causing its velocity to decrease rapidly and the curvature perturbation to grow on super-horizon scales. As a result, the primordial power spectrum can be amplified by several orders of magnitude over a narrow range of scales, leading to the collapse of overdense regions into PBHs after horizon re-entry in the radiation era [Garcia-Bellido:2017mdw, Motohashi:2017kbs, Germani:2017bcs, Byrnes:2018txb, Atal:2018neu, Figueroa:2021zah, Ragavendra:2023ret, Ballesteros:2024zdp, Ballesteros:2024pbe, Fujita:2025imc, Namjoo:2025hrr]. Such an enhancement of the curvature perturbation generically leads to non-Gaussianities (see [Pi:2024lsu] for a review), which can significantly alter the critical conditions for black hole formation, as shown by relativistic numerical studies [2020JCAP...05..022A, 2022JCAP...05..012E, Shimada:2024eec, Inui:2024fgk].
@@
The value of $\beta$ is related to the second derivative of the potential at the top of the small barrier. It was also realized that the relation is not invertible when the linear perturbation $\zeta_{G}$ exceeds $\beta^{-1}$. However, the probability for this to happen is finite, which hinted to a different channel for PBH formation. Indeed, while the barrier slows down the inflaton—enhancing curvature perturbations—it may also trap regions of the field that fluctuate backward, preventing them from crossing the barrier for $\zeta_{G}\gtrsim\beta^{-1}$. These trapped domains evolve into vacuum bubbles, providing an alternative channel of PBH formation distinct from the usual collapse of large adiabatic perturbations [2019JCAP...09..073A, 2020JCAP...05..022A, Escriva:2023uko]. Bubbles in this case form when the background field approaches the top of the barrier, leading to a sharply peaked distribution of bubble sizes. The precise form of this distribution was determined through dedicated numerical simulations [Escriva:2023uko], which found a scaling behavior for the bubble size. It was also shown that vacuum bubbles (surrounded by type-II fluctuations) can contribute higher PBH abundance than those forming from adiabatic perturbations, if the non-Gaussianity is large, i.e. $\beta\gtrsim 3.1$.
@@ 3 Statistics of initial conditions and shape dispersion
in $U_{+}$ and $U_{-}$ respectively. Here we have introduced $\sigma_{+}^{2}=\int_{U_{+}}{\cal P}_{\delta\phi}\ d\ln k\equiv\tilde{f}\sigma_{\delta\phi}^{2}$, and the complementary $\sigma_{-}^{2}=(1-\tilde{f})\sigma_{\delta\phi}^{2}$. This leads to
$\Delta W={\mu_{\Delta}^{2}\over\sigma_{\delta\phi}^{2}}{1\over\tilde{f}(1-\tilde{f})}$.
For fixed $\mu_{\Delta}$ the cost $\Delta W$ is minimized for $\tilde{f}=1/2$, when both subsets in the partition share equal weight in the variance of the field, i.e. $\sigma_{+}^{2}=\sigma_{-}^{2}$. In such case we have $\Delta W=4\mu_{\Delta}^{2}\sigma_{\delta\phi}^{-2}$. Then, from (3.13) the $n=1$ deviation from the mean corresponds to $\mu_{\Delta}=(1/2)\sigma_{\delta\phi}$. It follows from (3.10) and (3.14) that $|\Delta_{\phi}(r)|\leq\mu_{\Delta}$, and therefore we have
@@
Finally, to account for the likelihood of the shapes deviating from the mean profiles, we shall weigh those realizations by the factor
$\Upsilon(n,m)=\textrm{Erfc}(\mid n\mid/\sqrt{2})\textrm{Erfc}(\mid m\mid/\sqrt{2}).$ (3.31)
An example of the shapes can be found in Fig.5. There, we also display the anticorrelation between the two-point function $\Psi_{\phi}/\Psi_{\pi}$.
@@ 5 Quantum diffusion
Note, on the other hand, that here we considered the specific case of a potential with a flat plateau. For a plateau with some slope, the dynamics of vacuum-bubble formation can be significantly different, since the classical drift $v\equiv V^{\prime}/3H^{2}$ will make it harder for the inflaton to stay trapped. However, our picture should still apply in the regime where the slope of the plateau is sufficiently small, $v\ll H/2\pi$, so that classical drift is slow compared with diffusion. In that case, when the slope is towards the absorbing boundary, and for a reflecting boundary which is far away, $L\gg H^{2}/(2\pi v)\gg H$, the fractal dimension of the inflating regions is given by $D=3-2\pi^{2}v^{2}/H^{2}$, also very close to 3, while for $H\ll L\ll H^{2}/(2\pi v)$ we recover Eq.(5.6).
@@ 6 Adiabatic fluctuations and generalized $\zeta[\zeta_{G}]$
$\zeta[\zeta_{G}]=-\ln\left[\frac{\zeta_{G}A}{2}+\left(-\frac{q}{2}+\sqrt{\Delta}\right)^{1/3}+\left(-\frac{q}{2}-\sqrt{\Delta}\right)^{1/3}\right],$ (6.17)
@@
$q=-1+\zeta_{G}/\zeta^{\rm div}_{G}-A^{3}\zeta^{3}_{G}/4,$ (6.18)
$\Delta=\left[6(A+2)(\zeta_{G}-\zeta_{G}^{\rm div})\right]\left[6(A+2)(\zeta_{G}-\zeta_{G}^{\rm div})-2A^{3}\zeta^{3}_{G}\right]/64,$ (6.19)
@@
However, from Fig. 9 the value of $A\approx-0.1$ near the peak of the distribution is significant, and cannot be completely neglected. [^6] Here, we are mostly interested in the non-perturbative regime, relevant for PBH production. However, a perturbative expansion may be useful for studying stochastic backgrounds of induced gravitational waves. Expanding (6.21) in powers of $\zeta_{G}$, as $\zeta=\zeta_{G}+(3/5)f_{\mathrm{NL}}(k)\zeta_{G}^{2}+$... we find that the effective non-Gaussianity parameter $f_{\mathrm{NL}}(k)=(5/2)-(5/3)(k/k_{*})^{2}-(5/12)(k/k_{*})^{4}+$... runs with $k$. This behaviour is expected to hold in the regime $(k/k_{*})<0.5$, where the Pearson coefficient is maximal, $\gamma_{k}\approx-1$ (see Fig. 9). A qualitatively similar running of $f_{\mathrm{NL}}$, decreasing towards the UV, has been reported recently in [Namjoo:2025hrr]. For earlier works on this topic, see e.g. [Passaglia:2018ixg, Ragavendra:2020sop, Ozsoy:2021pws, Motohashi:2023syh].
@@
For each curvature profile $\zeta(r)$ generated with an amplitude $\mu<\mu^{\rm bub}_{c}$ (which therefore does not lead to the formation of vacuum bubbles), we perform relativistic numerical simulations in spherical symmetry to track the collapse of adiabatic curvature fluctuations, from super-horizon scales up to the moment of PBH formation. In particular, we use the numerical code SPriBHoS-II [Escriva:2025eqc] (see also [escriva_solo]), which employs a new methodology based on the Misner–Sharp formalism [PhysRev.136.B571] to follow the hydrodynamic collapse of generic curvature fluctuations (type-I and type-II) by using the trace of the extrinsic curvature. We refer the reader to that work for further details on the numerical procedure. We then iterate over different realizations of $\mu$ to determine the critical threshold conditions, defining this value as $\mu^{\rm adi}_{c}$. This can be converted into a threshold associated with the Gaussian curvature fluctuation $\zeta_{G}$, denoted by $\mu^{\zeta_{G}}_{c}=\mu^{\rm adi}_{c}/\dot{\phi}_{\rm bkg}(N_{*})$, from which we define the peak height $\nu^{\rm adi}_{c}=\mu^{\zeta_{G}}_{c}/\sigma_{\zeta_{G}}(N_{*})$, being $\sigma_{\zeta_{G}}(N_{*})$ the width of the power spectrum of $\mathcal{P}_{\zeta_{G}}(N_{*},k)$. The corresponding numerical results for these cases are shown in Table 1. For the mean profile $(n,m)=(0,0)$ we find that $\mu^{\rm adi}_{c}\approx 2.455\cdot 10^{-5}$ and $\mathcal{C}_{c}(r_{m})\approx 0.485$ (being $r_{m}$ the location of the maximum of the critical compaction function), and we also list the threshold for the other realizations $(n,m)$. When comparing the threshold obtained for the mean profile with that of the analytical template using the case A=0.1A=-0.1 (motivated by Fig.9, which is related to the maximum of the power spectrum),
In spite of the limitations of the analytic template near the divergent point, which we mentioned above, this still provides reasonably good agreement with the numerical $\delta N$ computation for lower values of $\zeta_{G}$, closer to the adiabatic threshold. For instance, for $(n,m)=(0,0)$ the full numerical simulation gives $\mu^{\zeta_{G}}_{c}=0.346$, while using the template Eq. (6.17) we obtain $\mu^{\zeta_{G}}_{c}\approx 0.334$, which is about $\sim 3\%$ lower. The difference is even smaller ($\sim 0.6\%$) when comparing with the peak value of the compaction function, $\mathcal{C}_{c}(r_{m})\approx 0.482$, using the same template. On the other hand, when taking $A=0$, corresponding to the template often used for USR Eq. (6.22), we obtain $\mu^{\zeta_{G}}_{c}\approx 0.317$ and $\mathcal{C}_{c}(r_{m})\approx 0.472$, with significantly larger deviations of $\sim 8\%$ and $\sim 3\%$, respectively, when compared with the numerical $\delta N$. We emphasize that an accurate determination of the PBH formation threshold, accounting for profile dependence, is crucial for reliably quantifying the resulting PBH abundance. In this work, such accuracy is achieved through dedicated numerical simulations.
@@
$\beta_{1}(n,m)$ $=\frac{\mathcal{I}(\nu^{\rm bub}_{c}(n,m),\infty)}{\mathcal{I}(\nu^{\rm bub}_{c}(0,0),\infty)}\frac{\Upsilon(n,m)}{\Upsilon(0,0)},$ (6.23)
$\beta_{2}(n,m)$ $=\frac{\mathcal{I}(\nu^{\rm adi}_{c}(n,m),\nu^{\rm bub}_{c}(n,m))}{\mathcal{I}(\nu^{\rm adi}_{c}(0,0),\nu^{\rm bub}_{c}(0,0))}\frac{\Upsilon(n,m)}{\Upsilon(0,0)},$ (6.24)
@@ Acknowledgments
A.E. thanks the support from the YLC program at the Institute for Advanced Research, Nagoya University. This work is supported in part by the National Key Research and Development Program of China Grant No. 2021YFC2203004. JG is supported by MCIN/AEI/PID2022-136224NB-C22, by the “Center of Excellence Maria de Maeztu 2025-2029” grant CEX2024-001451-M, funded by MICIU/AEI/10.13039/501100011033, and by 2021-SGR00872 funded by AGAUR. S.P. is supported by the National Natural Science Foundation of China (NSFC) Grants Nos. 12475066 and 12447101, and JSPS KAKENHI grant No. 24K00624. S.P. acknowledges the APCTP symposium APCTP-2025-IW1117 and the 41st annual IAP symposium “Inflation 2025” for the helpful feedback from colleagues.